# Multi-MeV electrons observed by CRRES during solar cycle 22

### v1

@@ 6 Summary & Conclusions
This study has utilised the CRRES-HEEF instrment to analyse relativistic electrons during solar cycle 22 with well-resolved measurements extending from 0.65 to 4.55 MeV. The slot region was observed to be flooded with electrons up to 4.55 MeV following three intense geomagnetic storms in the latter half of the mission. Comparable events have only been observed on a few occasions since, for example during and following the Hallowe’en storm period of 2003 in solar cycle 23 [Baker et al., 2013, Meredith et al., 2009] and during the Mother’s day/Gannon storm of May 2024 in solar cycle 25 [Pierrard et al., 2024], with solar cycle 24 identified as possessing an “impenetrable” barrier preventing high energy electrons from reaching below L*= 2.8 [Baker et al., 2014].
While the lack of continuous upstream solar wind observation in 1991 precludes further analysis of the upstream drivers of these slot-filling events observed by CRRES, the regularity of this phenomenon during the latter half of the mission can only be interpretted as presenting a persistent risk during large, but non-extreme, geomagnetic storms and therefore to satellites operating in slot region orbits. This directly affects constellations such as the O3B which operates at L=2.2.
The average electron loss timescales are compared to other solar cycles and several differences were highlighted. These include longer-lasting sub-MeV electrons near the inner edge of the outer belt, several slot-filling events which allowed the first definitive timescales at multi-MeV energies to be derived, and faster decaying fluxes beyond L* $\approx$ 5. These differences are associated with higher levels of geomagnetic activity. Several reasons were put forth to explain this which included higher levels of solar and geomagnetic activity during solar cycle 22 than in subsequent cycles.
@@
To examine the individual decay periods in more detail we also parameterise the results according to geomagnetic activity. This reveals several important trends. At higher L shells, the results are generally well-ordered by geomagnetic activity with high Kp decays often driving shorter loss timescales. Near the inner regions of the outer belt and in the slot region, however, this did not appear to be the case with a series of slow decays corresponding to high levels of activity. These persistent trends suggest that pitch angle diffusion cannot solely explain the evolution of fluxes during this periods, with other competing transport processes likely contributing.
@@
A further outcome of this investigation is that this HEEF dataset provided valuable data regarding radiation belt dynamics during solar cycle 22, either for stand-alone investigations or boundary conditions for radiation belt modelling. Care should be taken however, particularly when interpreting absolute fluxes, with sight of the underlying energy and pitch angle spectra.

### v2

@@ 7 Summary & Conclusions
This study has utilised the CRRES-HEEF instrument to analyse relativistic electrons during the maximum of solar cycle 22 with well-resolved measurements extending from 0.65 to 4.55 MeV. The slot region was observed to be flooded with electrons up to at least 4.55 MeV following two severe and one strong geomagnetic storms in the latter half of the mission. Comparable events have only been observed on a few occasions since, for example during and following the Hallowe’en storm period of 2003 in solar cycle 23 [Meredith et al., 2009, Baker et al., 2013] and during the Mother’s day/Gannon storm of May 2024 in solar cycle 25 [Pierrard et al., 2024, Li et al., 2025], with solar cycle 24 identified as possessing an “impenetrable” barrier preventing high energy electrons from reaching below L*= 2.8 [Baker et al., 2014]. The regularity of this phenomenon during the latter half of the mission can therefore only be interpreted as presenting the possibility of reoccurring during sustained periods of elevated solar activity, and therefore representing a direct risk to the increasing number of medium earth orbit (MEO) satellites operating in the slot region.
@@
To examine the individual decay periods in more detail we parameterise the results according to the wide spread in geomagnetic activity during this peiod. This reveals several important trends. At higher L shells, the results are generally well-ordered by geomagnetic activity with high Kp decays often driving shorter loss timescales. Near the inner regions of the outer belt and in the slot region, however, this did not appear to be the case with a series of slow decays corresponding to high levels of activity. These persistent trends suggest that pitch angle diffusion cannot solely explain the evolution of fluxes during this periods, with other competing transport processes, such as energy and radial diffusion, likely contributing. The large geomagnetic storms which flooded the slot region, were also discussed in terms of their severity, with Kp, Dst and aa indices often diverging. While the Kp index has been highlighted as better predicting radiation belt dynamics [Borovsky and Shprits, 2017], the lack of strict ordering of electron lifetimes by this index indicates a more sophisticated metric may be better for parametrising electron lifetimes.
The HEEF electron loss timescales were then compared to lifetimes from other solar cycles and several differences were highlighted. These include longer-lasting sub-MeV electrons near the inner edge of the outer belt, the several slot-filling events which allowed the first definitive timescales at multi-MeV energies to be derived, and faster decaying fluxes beyond L* $\approx$ 5. These differences are associated with higher levels of geomagnetic activity and match similarly elevated sub-MeV lifetimes in this region from solar cycle 23 and 20. This was therefore suggested to result from higher levels of solar and geomagnetic activity during the maximum of solar cycle 22 than VAP observations in cycle 24.
@@
A further outcome of this investigation is that this HEEF dataset provided valuable data regarding radiation belt dynamics during the maximum of solar cycle 22, either for stand-alone investigations or boundary conditions for radiation belt modelling. Care should be taken however, particularly when interpreting absolute fluxes, with sight of the underlying energy and pitch angle spectra.
@@ Acknowledgements
J.P. and S.R. acknowledge University of Warwick Undergraduate Research Support Scheme (URSS) bursaries. RTD acknowledges a Science and Technology Facilities Council Ernest Rutherford Fellowship ST/W004801/1 and UKSA grant ST/Y005635/1. NPM and SAG were supported in part by Natural Environment Research Council grants NE/V00249X/1 (Sat-Risk), NE/X000389/1 and NE/R016038/1. The views expressed are those of the authors and do not necessarily reflect the official policy or position of the Department of the Air Force, the Department of Defense, or the U.S. government. We thank the reviewers and editor for their helpful comments on this manuscript.
Open Research Section
The CRRES-HEEF dataset can be accessed on Zenodo at: Johnston et al. [2014a]. The theoretical diffusion coefficients, loss timescales and pitch angle distributions reported can be accessed on the NERC EDS UK Polar Data Centre at Glauert and Ross [2024].